If the sum of four interior angles of a hexagon is 480⁰, what is the sum of the remaining 2 angles
Answer by math_tutor2020(3817)


S = sum of the interior angles of a polygon with n sides
S = 180(n-2)
S = 180(6-2)
S = 720

All hexagons have their six interior angles add to 720 degrees.

480 degrees is already accounted for, so we have 720 - 480 = 240 degrees left over (for the remaining two angles that add to this).

Answer: 240 degrees